# Eye-tracked Virtual Reality: A Comprehensive Survey on Methods and Privacy Challenges

### v2

@@ IV-C1 Human Cognition
Unlike the tangible fields, the realm of human cognition is characterized by abstraction and is hence more elusive. To close this gap, eye-tracked VR has shown great potential in studying human cognition. Gaze features such as eye movement patterns, pupil diameters, and aggregated eye features such as fixation and saccade durations have been widely used as indicators for human cognitive processes. This section of our paper summarizes the research on the relationship between eye movements and cognition.
@@
Bozkir et al. [230] studied cognitive load in driving simulations in VR and showed that with machine learning, it is possible to predict human cognitive load with an accuracy of over 80% by incorporating eye features, particularly by using pupil diameters. Souchet et al. [231] conducted cognitive load prediction using features from different modalities, including gaze features, electrocardiogram, and electrodermal activity. The authors found that electrocardiogram features were the most relevant features correlated with self-reported subjective cognitive load levels compared to the other modalities in their setup. Furthermore, Duchowski et al. [232] pointed out the drawbacks of pupillometry in cognitive load assessment that are caused by ambient illumination and off-axis distortion and suggested the wavelet-based estimation of pupil diameter oscillation frequency as an alternative feature. The proposed metric showed great potential to classify task difficulty in a user study where 17 participants did math calculations.
@@
Researchers also utilized cognitive load prediction tasks in the aviation field. Luong et al. [233] utilized features from various modalities, including pupil features such as pupil diameter size, pupil diameter amplitude, and constriction and dilatation speed of the pupil to evaluate the cognitive load in a VR flight simulation. Using random forests, the authors could classify four levels of cognitive load with an accuracy of up to 65%. Furthermore, Wilson et al. [234] studied cognitive load in flight simulation using various sensing modalities with a deep learning model. The authors fed pupil dilation and blink features into their models and obtained 80% accuracy in cognitive load estimation.
@@
Bækgaard et al. [235] examined the correlation between pupil dilation and the difficulty index in Fitts’ Law [236]. Fitts’ Law describes the relationship between the time, distance, and accuracy of an individual’s movement in reaching a given target while considering the level of cognitive load or task difficulty, enabling the assessment of task difficulty or cognitive load by defining the difficulty index as a quantitative measure. Findings of Bækgaard et al. indicate that tasks that demand higher cognitive load lead to a larger pupil diameter. In contrast, there is no significant relationship between pupil diameter and the Fitts’ Law index in motor task complexity.
@@
Measuring pupil dilation is challenging due to its sensitivity to illumination [237]. To overcome this, John et al. [238] attempted to exclude the impact of the light source on pupil dilation in 2D monitors and HMDs using three different models. A linear calibration model yielded prediction errors of 0.42 and 0.60 mm for pupil constriction and dilation in 2D, and an exponential model had an error of 0.27 mm in VR. Similarly, Eckert et al. [237] presented a method to correct for light-induced pupil diameter changes, and found that estimating luminance by using the weighted average of the fixation and background area provides the best performance. More related to applications, Gao et al. [239] recently compared the effects of five VR locomotion methods on cognitive response, including arm swinging, dash, grappling, joystick, and teleportation. The authors employed gaze features, such as blink rates and fixation durations, to measure cognitive load and discovered that locomotion methods could significantly influence the cognitive states in VR, while joystick and teleportation resulted in lower cognitive load in participants.
@@
Human emotions such as fear and frustration also play an important role in the human cognitive state, and an increasing number of studies have been devoted to identifying the relationship between gaze behavior and human emotion in virtual spaces. For instance, Pan et al. [240] used pupil diameter changes to measure the influence of the fear of the sea on working memory. The authors designed n-back memory tests for 29 subjects and found that pupil diameter changes noticeably increased in the underwater simulation in all tasks. Luong and Holz [241] investigated users’ pupillometry responses along with pulmonary, electrodermal, and cardiac data to capture sensations of fear, frustration, and insight. Insight refers to the moment when a player suddenly understands the solution to a puzzle or clue. The authors achieved an F1 score of over 71% to predict these senses using classification models, including logistic regression and support vector machine models. Furthermore, Stoeve et al. [242] used machine learning by utilizing gaze, pupil diameter, and blink information to successfully estimate goalkeeper stress in VR, which is relatable to cognitive load and processes.
@@
Apart from estimating the cognitive load, inferences about human cognition have also been used as a supportive source of information in various application domains, especially to enhance the interaction experience in VR. Lindlbauer et al. [243] utilized cognitive load, which was estimated by the index of pupil activity metric based on pupil dilation change, to optimize and eliminate the manual adjustment process of the level of information displayed in mixed reality glasses. The authors showed that their approach reduced the number of secondary task interactions by 36%. Kübler et al. [244] utilized cognitive load as an indicative factor to detect hazard situations in a driving simulator. SVM models trained with wavelet transformation components extracted from pupil diameter data were used to detect hazardous situations. The authors demonstrated that pupil dilation could help indicate hazard perception in driving simulations. However, they also noted that it cannot be relied upon as a stand-alone detection mechanism due to the number of false positives.
@@
When we consider applications of human cognition, one of the most important domains is medicine, as human cognitive processes are related to neurological and psychological diseases. In addition, since it is possible to manipulate and control experimental conditions very precisely in VR, researchers often employ VR and eye tracking together in this domain. For instance, to understand the response of children with autism spectrum disorder (ASD) to facial expressions, Bekele et al. [245] conducted a user study with 10 ASD patients and 10 typically developing (TD) peers in multiple emotion recognition tasks using a facial expression system in VR. By analyzing participants’ gaze features such as ROIs, fixation durations, blink rates, and pupil diameters, the latter being indicative of cognitive load, the authors found that the way individuals with ASD identify emotional faces is significantly different from their TD peers.
@@
In another work, Bekele et al. [246] introduced a novel multi-modal adaptive social interaction in a VR platform and suggested using gaze information for adaptive intervention. In particular, the authors used ROIs to adjust the face occlusion of virtual characters in real-time. In their emotion recognition experiments with six ASD and six age-matched children, a performance increase of 3% associated with the novel gaze-sensitive mechanism was observed. Similarly, Kim et al. [247] presented a VR-based interactive social skill training system (VISTA) that uses eye-tracking data, including ROIs and pupil features, to understand the characteristics of people with ASD. In a user study of 20 participants, the authors found that the ASD group was highly engaged with VISTA while showing a dramatic difference from the neurotypical group in biological signals, such as a larger variation in pupil diameters, which generally indicates a higher cognitive load.
@@
The research in the medical domain that includes human cognition, eye tracking, and VR is not limited to the ASD research. Orlosky et al. [248] presented a remote diagnosis system in VR for neurodegenerative diseases like Parkinson’s. The authors designed various tasks to elicit abnormal eye behaviors such as ocular tremors, square wave jerks, and abnormal pursuits, which could indicate neurodegenerative diseases and cognitive load. Based on the experimental results, the authors claimed that the VR interface successfully elicited five types of abnormal eye movements. In contrast, eye physicians could identify three out of four abnormal eye behaviors. The authors also implemented three different visualization techniques for eye movements, which can be helpful for doctors, as humans can hardly identify small movements. While the diagnosis of such diseases could not only be made with the help of eye tracking and VR, because those diseases are far more complex for only machines and data-driven practices to identify, such systems can support doctors in their regular workflows. Table VI provides an overview of papers discussed in Section IV-C1, including their purposes, eye-tracking features, and the number of participants involved, whereas Table V includes abbreviations for the features.
@@
TABLE VI: An overview of the papers related to human cognition.
Paper # Participants Purpose Eye-tracking Features [230] 16 Cognitive load prediction in driving simulation pup [231] 92 Cognitive load prediction pup, sac, blnk [232] 17 Cognitive load prediction pup, freq [233] 75 Cognitive load prediction in flight simulation pup [234] 40 Cognitive load prediction in flight simulation pup, blnk [235] 27 Cognitive load analysis in physiology pup [238] 24 Cognitive load analysis pup [237] 17 Cognitive load analysis pup [239] 15 Cognitive response analysis for VR locomotion blnk, fix, sac [240] 29 Emotion recognition pup [241] 24 Emotion recognition pup [242] 30 Stress classification gaze, pup, blnk [243] 12 Cognitive load prediction pup [244] 31 Cognitive load analysis for hazard detection pup [245] 20 Understanding ASD gaze, fix, blnk, pup [246] 12 Understanding ASD for adaptive intervention gaze, pup, blnk, fix, sac [247] 20 Understanding ASD for social skill training gaze, pup [248] 16 Parkinson’s disease detection gaze, pup
@@ IV-C2 Visual Attention
This section gives a holistic view of prior works linking eye-tracked VR and human visual attention. Human visual attention is related to where people look at the scene, mainly with their eye gaze, and often correlates with cognitive and perceptual abilities. In addition, visual attention is affected by the 3D scene design and stimulus-related factors in VR. To understand these, researchers have studied various aspects of human visual attention.
@@
Schmitz et al. [249] explored how peripheral flicker and central arrow stimuli differently impacted visual attention guidance in virtual panoramic videos. Based on their analysis with 25 participants, the researchers reported that the participants preferred the central arrow stimuli over the peripheral flicker stimuli, perceiving the former as more rewarding. Additionally, the authors claimed that traditional attention mechanisms might not fully apply to panoramic videos. Soret et al. [250] investigated the effect of endogenous and exogenous cues to orient visual attention with eye tracking and found that such cue types reduce the users’ reaction time as initially anticipated. Liu et al. [251] performed unordered tasks more efficiently over time by employing dynamic visual cues to guide users’ visual attention towards specific substeps of the performed task. This method allowed users to determine the sequence of task completion, which subsequently influences the presented set of cues, which are dynamically updated based on hand proximity and eye gaze. The authors stated that visual cues based on eye gaze improve task completion times more significantly than cues based on hand proximity.
Lange et al. introduced HiveFive [252] as an attention guidance technique designed to direct the users’ attention to context-relevant points using swarm motion. The proposed technique achieves lower response latency, resulting in a less adverse impact on immersion. Researchers also proposed visual attention guidance in VR-based driving simulators. Bozkir et al. [253] provided gaze-aware warning cues for pedestrian crossings in a low-cost VR to enhance driver attention. The experimental results, which include metrics such as nearest distances to pedestrians, participant pupil diameters, and drivers’ inputs on pedals, showed that eye-tracked VR might be an effective tool for driver training in safety-critical situations due to the effectiveness of the visual cues.
@@
Notifications are often considered a form of external stimulus that may interrupt the user during virtual experiences. Gaze-based visual attention analysis could help to understand the impact of notifications and potentially provide a less disruptive VR experience. To this end, Hsieh et al. [254] explored the perception of notifications by VR users when they received them in a message form from the real world. The authors provided design suggestions to eliminate the disruptive effect of these notifications, such as by measuring users’ level of engagement using eye-tracking data to find the convenient times for notifications and locating the message according to the user’s peripheral area. Similarly, Chen et al. [255] proposed a technique using a deep learning model by processing time series sensor data, including gaze angles and gaze shift rates, to predict the optimal time for notification. The authors’ deep learning model predicted a favorable time for notifications with 71% precision, and further improvements were observed when the information related to user activities and engagement. Similarly to the aforementioned works about notifications, the assessment of visual attention is considered useful for identifying the appropriate time and region to insert unperceived modifications within a virtual scene. Mise-Unseen [256] by Marwecki et al. is an approach that injects scene changes imperceptibly into the users’ field of view. To adjust the injection time, the authors used eye-tracking data to assess the user’s attention, intent, and spatial memory. The authors demonstrated that using gaze data in combination with masking techniques enables the insertion of modifications without being perceived by the user.
@@
Visual attention analysis is useful for the design phase of virtual spaces. 3D virtual spaces could be designed to support the primary task objectives by analyzing and understanding the visual attention. To this end, Alghofaili et al. [257] presented an attention-based approach to optimize the placement of visual elements and design a 3D environment considering predefined goals. The authors employed a regression model to predict gaze duration and showed an optimization efficiency, combining cost functions for regularization, number of elements, and primary design objective related to the visual elements to be placed. Hillaire et al. [258] introduced an algorithm for exploring virtual scenes by simulating human visual attention. The proposed algorithm, which can also be applied during first-person exploration that includes walking and turning movements, uses a surface element-based representation instead of a mesh-based representation. The authors stated that the proposed algorithm and model could be used to optimize the quality of the scene to accelerate the rendering process.
@@
With VR, it is easier to assess the joint attention of users, as with each HMD and user, it is possible to get precise gaze and head orientations in a virtual space. As the information on visual attention in VR also promotes interaction and collaboration, joint attention has been a focal point for some. Špakov et al. [259] analyzed the effect of eye-gaze- and head-gaze-based visual attention sharing in cooperative games. In such VR games, sharing eye-gaze led to higher subjective ratings of teamwork and shorter game duration than sharing head-gaze. Similarly, Kasahara et al. [260] evaluated the effect of the shared egocentric videos with gaze locations in perspective sharing. The behavior changes and decisions of the participants with experience sharing were analyzed while sharing parallel views. During the drawing activity, the researchers observed that individuals can develop behaviors to complement their partner’s memory and decision-making. Furthermore, even in complex situations, individuals can create mechanisms to comprehend their physical embodiment and spatial relationship with their peers.
In the realm of medicine, impairments in joint attention can serve as an early indicator of ASD. Training and enhancing the joint attention mechanisms are essential in interventions for individuals, for instance, with ASD. Mei et al. [261] attempted to address skills training for the joint attention of ASD individuals using customizable virtual humans (CVH). The researchers designed an educational drum-playing scenario in an eye-tracked virtual environment for 10 ASD individuals and measured visual attention with the help of gaze data. The authors’ user study showed that CVH improved participants’ visual focus on relevant regions of the scene at the cost of increased reaction time.
@@
Eye tracking has also found applications in pedagogy, particularly in virtual learning spaces. Especially with the impact of the COVID-19 pandemic, remote and immersive learning spaces have become popular, especially in the form of virtual classrooms. In such simulations and setups, the visual attention of students and teachers plays an important role in the learning processes. To facilitate this, Ahuja et al. [262] proposed a novel 3D digital twin classroom simulation that could be accessed through both VR headsets and a web interface and incorporated a computer vision-based head- and eye-tracking system. Two cameras were deployed for the system to track students and teachers, respectively, and gaze data were visualized in diverse forms, such as students’ saliency map on the board and intersections of instructors’ gaze on student planes. A controlled study showed that this novel non-HMD classroom gaze system almost halved the gaze prediction error of its predecessors. In contrast, a reliability of 92.54% was reported in an in-the-wild assessment.
@@
Focusing more on the virtual simulations, Gao et al. [263] utilized several gaze features such as fixation durations, saccade durations, and pupil diameters to investigate the influence of various classroom manipulations on learners in a VR classroom, including different sitting positions, virtual avatar styles, and peer hand-raising behaviors. The authors found that such manipulations affect user attention and cognition differently, which could lead to different learning and engagement outcomes in the long term. Bozkir et al. [264] also studied the same manipulations in an eye-tracked VR classroom. In this work, the authors focused on virtual objects and attention times on the objects related to learning and engagement in the classroom. The authors showed that learners sitting in the front part of the virtual classroom paid more attention to the teacher and board, whereas others who sat in the back visually focused more on their peers. The authors also discovered that cartoon- and realistic-styled avatars grabbed different amounts of attention during the classroom discourse.
Later, Gao et al. [265] explored various gaze features in a virtual lecture to scrutinize the effects of social interactions. The authors considered peer-learner hand-raisings as a cue of social interaction in their virtual classroom. They found that hand-raising animations and the number of peer learners who raise their hands affect students’ attention in different ways. Attention and cognition analyses toward such cues in virtual learning spaces are especially important as student self-concepts could be influenced positively or negatively by the same amount of peers raising their hands constantly. Using the same VR classroom space, Gao et al. [266] also showed that boys and girls visually behave differently with their eye gaze in the classroom discourse. The authors indicated that utilization of this information could be of help for personalized learning support, and their explainable machine learning model resulted in a gender classification accuracy of over 70%.
@@
Complementary to the aforementioned research, Rahman et al. [267] focused on the perspective of the teacher during a lecture in VR and proposed six different gaze visualization techniques from the teachers’ view to identify distracted students. These visualization techniques covered gaze ring, gaze disk, gaze arrow, gaze trail, gaze trail with arrows, and gaze heatmap. The user study results showed that gaze trail visualization outperformed the others in terms of popularity, and the application of 3D gaze heatmaps was problematic. In addition, Khokhar and Borst [268] found in an educational VR setup that users perceive increased interactivity of the teacher agents as more natural and appropriate, providing insights on how to design virtual interaction experiences in learning spaces. Similar to papers about human cognition, we summarize the papers related to visual attention in Table VII by including the stimulus types associated with each paper.
@@
TABLE VII: Overview of the papers related to human visual attention.
Paper # Participants Purpose Eye-tracking Features Visual Stimuli [249] 25 Attention analysis in panaromic videos gaze Cues [250] 20 Attention guidance in sandwich preparation gaze Cues [251] 15 Attention guidance in unordered tasks gaze Dynamic cues [252] 20 Attention guidance gaze Swarm motion [253] 16 Attention guidance in driving simulation pup Cues [254] 40 Attention analysis gaze Notifications [255] 20 Attention analysis gaze Notifications [256] 15 Attention analysis for scene changes gaze Scene changes [257] 23 Attention analysis for environment design gaze N/A [258] 12 Attention analysis for environment design gaze N/A [259] 40 Joint attention analysis in collaboration gaze Shared gaze [260] 40 Joint attention analysis in collaboration gaze Shared gaze [261] 10 Joint attention training for patients with ASD gaze N/A [262] 13 Attention analysis in education saliency, gaze N/A [263] 288 Attention analysis in education fix, sac, pup N/A [264] 280 Attention analysis in education gaze N/A [265] 280 Attention analysis in education pup, fix, sac, gaze N/A [266] 280 Gender analysis in education pup, fix, sac, gaze N/A [267] 26 Gaze visualization in education gaze Visualized gaze

### v3

@@ IV-C1 Human Cognition
Unlike the tangible fields, the realm of human cognition is characterized by abstraction and is hence more elusive. To close this gap, eye-tracked VR has shown great potential in studying human cognition. Gaze features, such as eye movement patterns and pupil diameters, have been widely used as human cognitive process indicators.
@@
To this end, cognitive load prediction has been an important focus. [202] studied cognitive load in driving simulations in VR and showed that it is possible to predict human cognitive load with an accuracy of over 80% by incorporating eye features, particularly by using pupil diameters, with machine learning. [203] conducted cognitive load prediction using features from different modalities, including gaze features, electrocardiogram, and electrodermal activity. The electrocardiogram features were the most relevant features correlated with self-reported subjective cognitive load levels compared to the other modalities. Furthermore, [204] pointed out the drawbacks of pupillometry in cognitive load assessment that are caused by ambient illumination and off-axis distortion and suggested the wavelet-based estimation of pupil diameter oscillation frequency as an alternative feature. The proposed metric demonstrated potential for classifying task difficulty while participants were performing math calculations.
@@
Researchers also utilized cognitive load prediction tasks in the aviation field. [205] utilized features including pupil diameter size, pupil diameter amplitude, and constriction and dilatation speed of the pupil to evaluate the cognitive load in a VR flight simulation. Using random forests, the authors could classify four levels of cognitive load with an accuracy of up to 65%. Furthermore, [206] investigated cognitive load estimation in flight simulations, achieving 80% accuracy using pupil dilation and blink features.
@@
Studying the correlation between pupil dilation and the difficulty index in Fitts’ Law [207], which describes the relationship between the time, distance, and accuracy of an individual’s movement in reaching a given target while considering the level of cognitive load or task difficulty, [208] indicated that tasks that demand higher cognitive load lead to a larger pupil diameter. In contrast, there was no significant relationship between pupil diameter and the Fitts’ Law index in motor task complexity.
@@
Measuring pupil dilation is challenging due to its sensitivity to illumination [209]. To overcome this, [210] attempted to exclude the impact of the light source on pupil dilation in 2D monitors and HMDs using three different models. A linear calibration model yielded prediction errors of 0.42 and 0.60 mm for pupil constriction and dilation in 2D, and an exponential model had an error of 0.27 mm in VR. Similarly, [209] presented a method to correct for light-induced pupil diameter changes, and found that estimating luminance by using the weighted average of the fixation and background area provides the best performance. Related to its applications, [211] compared the effects of five VR locomotion methods on cognitive response, including arm swinging, dash, grappling, joystick, and teleportation, by employing gaze features, such as blink rates and fixation durations, to measure cognitive load and discovered that locomotion methods could significantly influence the cognitive states in VR.
@@
Human emotions, such as fear and frustration, also play a crucial role in the human cognitive state, and an increasing number of studies have been devoted to understanding the relationship between gaze behavior and emotions in VR. For instance, [212] used pupil diameter changes to measure the influence of the fear of the sea on working memory by designing n-back memory tests and found that pupil diameter changes noticeably in the underwater simulations. [213] investigated users’ pupillometry responses along with pulmonary, electrodermal, and cardiac data to capture sensations of fear, frustration, and insight, where the latter refers to the moment when a user suddenly understands the solution to a puzzle or clue. The authors achieved an F1 score of over 71% to predict these senses using classification models, including logistic regression and SVMs. Furthermore, [214] employed machine learning by utilizing gaze, pupil diameter, and blink information to successfully estimate goalkeeper stress in VR, which is relatable to cognitive load and processes.
@@
Apart from estimating the cognitive load, inferences about human cognition have also been used as a supportive source of information in various application domains, especially to enhance the interaction experience in VR. [215] utilized cognitive load, which was estimated by the index of pupil activity metric based on pupil dilations, to optimize and eliminate the manual adjustment process of the level of information displayed in MR glasses. This approach reduced the number of secondary task interactions by 36%. [216] utilized cognitive load as an indicative factor to detect hazard situations in a driving simulator. SVMs trained with wavelet transformation components extracted from pupil diameter data were used to detect hazardous situations. The results demonstrated that pupil dilation could help indicate hazard perception in driving simulations; however, it cannot be relied upon as a stand-alone detection mechanism due to the number of false positives.
@@
One of the most critical domains when considering applications of human cognition is medicine, as human cognitive processes are closely related to neurological and psychological diseases. In addition, since it is possible to manipulate and control experimental conditions very precisely in VR, researchers often employ VR and eye tracking together in this domain. For instance, to understand the response of children with autism spectrum disorder (ASD) to facial expressions, [217] conducted a user study and analyzed gaze features such as ROIs, fixation durations, blink rates, and pupil diameters, the latter being indicative of cognitive load, and found that the way individuals with ASD identify emotional faces is significantly different from their typically developing peers.
@@
In another work, [218] introduced a novel multi-modal adaptive social interaction in VR and suggested using gaze information for adaptive intervention. The authors used ROIs to adjust the face occlusion of virtual characters in real-time. In their emotion recognition experiments in the context of ASD, a performance increase of 3% associated with the novel gaze-sensitive mechanism was observed. Similarly, [219] presented a VR-based interactive social skill training system (VISTA) that uses eye tracking, including ROIs and pupil features, to understand the characteristics of people with ASD. The authors found that the ASD group was highly engaged with VISTA while showing a dramatic difference from the neurotypical group in biological signals, such as a larger variation in pupil diameters, which indicates a higher cognitive load.
@@
Research in the medical domain, which includes human cognition, eye tracking, and VR, is not limited to the ASD context. [220] presented a remote diagnosis system in VR for neurodegenerative diseases like Parkinson’s. The authors designed various tasks to elicit abnormal eye behaviors such as ocular tremors, square wave jerks, and abnormal pursuits, which could indicate neurodegenerative diseases and cognitive load. The VR interface successfully elicited five types of abnormal eye movements, whereas eye physicians could identify three out of four abnormal eye behaviors. The authors also implemented different visualization techniques for eye movements, which can be helpful for doctors, as humans can hardly identify small movements. While the diagnosis of such diseases cannot be made solely with the help of eye tracking and VR, as these diseases are far more complex for machines and data-driven practices to identify, such systems can support doctors in their regular workflows. Table VI provides an overview of papers in Section IV-C1, including their characteristics, eye-tracking features, and the number of participants, whereas Table V includes abbreviations for the features.
@@
TABLE VI: An overview of the papers related to human cognition. “#P” corresponds to the number of participants.
Characteristics #P Features
[202] Cognitive load prediction in driving 16 pup
[203] Cognitive load prediction 92 pup, sac, blnk
[204] Cognitive load prediction 17 pup, freq
[205] Cognitive load prediction in aviation 75 pup
[206] Cognitive load prediction in aviation 40 pup, blnk
[208] Cognitive load analysis in physiology 27 pup
[210] Cognitive load analysis 24 pup
[209] Cognitive load analysis 17 pup
[211] Cognitive responses during locomotion 15 blnk, fix, sac
[212] Emotion recognition 29 pup
[213] Emotion recognition 24 pup
[214] Stress classification 30 gaze, pup, blnk
[215] Cognitive load prediction 12 pup
[216] Cognition for hazard perception 31 pup
[217] Understanding ASD 20 gaze, fix, blnk, pup
[218] Adaptive intervention for ASD 12 gaze, pup, blnk, fix, sac
[219] Social skill training for ASD 20 gaze, pup
[220] Parkinson’s disease detection 16 gaze, pup
@@ IV-C2 Visual Attention
This section provides a comprehensive overview of prior works that link eye-tracked VR and human visual attention. Human visual attention is related to where people look in a scene, primarily through their eye gaze, and often correlates with cognitive and perceptual abilities. Additionally, visual attention is influenced by the design of 3D scenes and stimulus-related factors in VR.
@@
To understand these concepts, researchers have studied various aspects of human visual attention. [221] explored how peripheral flicker and central arrow stimuli differently impacted visual attention guidance in virtual panoramic videos. The authors reported that the participants preferred the central arrow stimuli over the peripheral flicker stimuli, perceiving the former as more rewarding. Additionally, they claimed that traditional attention mechanisms might not fully apply to panoramic videos. [222] investigated the effect of endogenous and exogenous cues to orient visual attention with eye tracking and found that such cue types reduce the users’ reaction time as initially anticipated. [223] performed unordered tasks more efficiently over time by employing dynamic visual cues to guide users’ visual attention towards specific substeps of the performed task. This method enabled users to determine the sequence of task completion, which subsequently influenced the presented set of cues, which were dynamically updated based on hand proximity and eye gaze. The authors stated that visual cues based on eye gaze improve task completion times more significantly than cues based on hand proximity. Studying attention guidance, [224] introduced HiveFive, a guidance technique to direct the users’ attention to context-relevant points using swarm motion, which achieves lower response latency, resulting in a less adverse impact on immersion.
@@
Notifications are often considered a form of external stimulus that may interrupt the user during virtual experiences, and visual attention analysis could help to understand the impact of notifications and potentially provide a less disruptive VR experience. To this end, [225] explored the perception of notifications by VR users when they received them in a message form from the real world. The authors provided design suggestions to mitigate the disruptive effect of these notifications, such as by measuring users’ level of engagement using eye tracking to identify convenient times for notifications and positioning the message in the user’s peripheral area. Similarly, [226] proposed a technique that utilizes a deep learning model to process time series sensor data, including gaze angles and gaze shift rates, for predicting the optimal time for notification. The authors’ deep learning model predicted a favorable time for notifications with 71% precision, and further improvements were observed in relation to user activities and engagement. The assessment of visual attention is also considered useful for identifying the appropriate time and region to insert unperceived modifications within a virtual scene. Mise-Unseen [227] is an approach that injects scene changes imperceptibly into the users’ field of view. To adjust the injection time, the authors used eye tracking to assess the user’s attention, intent, and spatial memory, demonstrating that gaze data in combination with masking techniques enables the insertion of modifications without being perceived by the user.
@@
Visual attention analysis is useful for the design phase of virtual spaces. 3D virtual spaces could be designed to support the primary task objectives by analyzing and understanding visual attention. To this end, [228] presented an attention-based approach to optimize the placement of visual elements and designed a 3D environment considering predefined goals. The authors employed a regression model to predict gaze duration and showed an optimization efficiency, combining cost functions for regularization, number of elements, and the primary design objective related to the visual elements to be placed. [229] introduced an algorithm for exploring virtual scenes by simulating human visual attention. The proposed algorithm, which can also be applied during first-person exploration that involves walking and turning movements, utilizes a surface element-based representation instead of a mesh-based representation. This approach enables the optimization of scene quality to accelerate the rendering process.
@@
With VR, it is easier to assess joint attention of users, as with each HMD and user, it is possible to get precise gaze and head orientations in VR. As the information on visual attention also promotes interaction and collaboration, joint attention has been a focal point for some. [230] analyzed the effect of eye-gaze- and head-gaze-based visual attention sharing in cooperative games. In such VR games, sharing eye-gaze led to higher subjective ratings of teamwork and shorter game duration than sharing head-gaze. Similarly, [231] evaluated the effect of the shared egocentric videos with gaze locations in perspective sharing. The behavior changes and decisions of the participants who shared experiences were analyzed, considering parallel views. During the drawing activity, the researchers observed that individuals can develop behaviors to complement their partner’s memory and decision-making. Furthermore, even in complex situations, individuals can create mechanisms to comprehend their physical embodiment and spatial relationship with their peers.
In the realm of medicine, impairments in joint attention can serve as an early indicator of ASD. Training and enhancing the joint attention mechanisms are essential in interventions for individuals, for instance, with ASD. [232] attempted to address skills training for the joint attention of ASD individuals using customizable virtual humans (CVH). In an educational drum-playing scenario, evaluations revealed that the CVH enhanced participants’ visual focus on relevant regions of the scene, albeit at the expense of increased reaction time.
@@
Eye tracking has also found applications in pedagogy, particularly in virtual learning spaces. Especially with the impact of the COVID-19 pandemic, remote and immersive learning spaces have gained popularity, particularly in the form of virtual classrooms. In such settings, the visual attention of students and teachers plays a crucial role in the learning processes. To facilitate this, [233] proposed a novel 3D digital twin classroom simulation that can be accessed through both VR HMDs and a web interface, incorporating a computer vision-based head- and eye-tracking system. Two cameras were deployed to track students and teachers, respectively, and gaze data were visualized in various forms, such as students’ saliency maps on the board and intersections of instructors’ gaze on student planes. A controlled study showed that this novel non-HMD classroom gaze system almost halved the gaze prediction error of its predecessors. In contrast, a reliability of 92.54% was reported in an in-the-wild assessment.
@@
Focusing more on the virtual simulations, [234] utilized several gaze features such as fixation durations, saccade durations, and pupil diameters to investigate the influence of various classroom manipulations on learners in VR, including different sitting positions, virtual avatar styles, and peer hand-raising behaviors. The authors found that such manipulations affect user attention and cognition differently, which could lead to different learning and engagement outcomes. [235] also studied the same manipulations in an eye-tracked VR classroom, focusing on virtual objects and attention times related to learning and engagement in the classroom. The authors demonstrated that learners seated in the front part of the virtual classroom paid more attention to the teacher and the board, whereas those seated in the back visually focused more on their peers. The authors also discovered that cartoon- and realistically styled avatars attracted different amounts of attention. Later, [236] explored various gaze features in a virtual lecture to scrutinize the effects of social interactions. The authors considered peer-learner hand-raisings as a cue of social interaction in VR and found that hand-raising animations and the number of peer learners who raise their hands affect students’ attention in different ways. Analyses of attention and cognition toward such cues in virtual learning spaces are critical, as student self-concepts can be positively or negatively influenced by these behaviors.
@@
Complementary to the research above, [237] focused on the teacher’s perspective during a lecture in VR and proposed six different gaze visualization techniques from the teacher’s view to identify distracted students. These visualization techniques covered gaze ring, gaze disk, gaze arrow, gaze trail, gaze trail with arrows, and gaze heatmap, and the results showed that gaze trail visualizations outperformed the others in terms of popularity, and the application of 3D gaze heatmaps was problematic. In addition, [238] found in an educational VR setup that users perceive increased interactivity of the teacher agents as more natural and appropriate, providing insights on the design implications for virtual learning spaces. Table VII summarizes the papers related to visual attention.
@@
TABLE VII: Overview of the papers related to human visual attention. “#P” corresponds to the number of participants.
Characteristics #P Features
[221] Attention analysis in panaromic videos 25 gaze
[222] Attention guidance in sandwich preparation 20 gaze
[223] Attention guidance in unordered tasks 15 gaze
[224] Attention guidance 20 gaze
[225] Attention analysis 40 gaze
[226] Attention analysis 20 gaze
[227] Attention analysis for scene changes 15 gaze
[228] Attention analysis for environment design 23 gaze
[229] Attention analysis for environment design 12 gaze
[230] Joint attention analysis in collaboration 40 gaze
[231] Joint attention analysis in collaboration 40 gaze
[232] Joint attention training for ASD 10 gaze
[233] Attention analysis in education 13 saliency, gaze
[234] Attention analysis in education 288 fix, sac, pup
[235] Attention analysis in education 280 gaze
[236] Attention analysis in education 280 pup, fix, sac, gaze
[237] Gaze visualization in education 26 gaze